How does the average pressure of an elephant standing on the ground compare to that exerted by a person standing on the ground? The person weighs 640 N. Each shoe has an area of about 0.016 m2 in contact with the floor. The elephant weighs 4.13×104 N and has four feet, each with an area of about 0.14 m2 in contact with the ground.




1
Expert's answer
2021-05-19T17:03:22-0400
"P_{el}=\dfrac{F}{A}=\dfrac{W_{el}}{A},""P_{el}=\dfrac{4.13\cdot10^4\ N}{4\cdot0.14\ m^2}=7.4\cdot10^4\ Pa,""P_{person}=\dfrac{640\ N}{2\cdot0.016\ m^2}=2\cdot10^4\ Pa,""\dfrac{P_{el}}{P_{person}}=\dfrac{7.4\cdot10^4\ Pa}{2\cdot10^4\ Pa}=3.7"

The average pressure of an elephant standing on the ground is 3.7 times greater than that eerted by the person.